# Not All Data are Good Labels: On the Self-supervised Labeling for Time Series Forecasting

### v2

@@ Abstract
Time Series Forecasting (TSF) is a crucial task in various domains, yet existing TSF models rely heavily on high-quality data and insufficiently exploit all available data. This paper explores a novel self-supervised approach to re-label time series datasets by inherently constructing candidate datasets. During the optimization of a simple reconstruction network, intermediates are used as pseudo labels in a self-supervised paradigm, improving generalization for any predictor. We introduce the Self-Correction with Adaptive Mask (Scam), which discards overfitted components and selectively replaces them with pseudo labels generated from reconstructions. Additionally, we incorporate Spectral Norm Regularization (snr) to further suppress overfitting from a loss landscape perspective. Our experiments on eleven real-world datasets demonstrate that Scam consistently improves the performance of various backbone models. This work offers a new perspective on constructing datasets and enhancing the generalization of TSF models through self-supervised learning. The code is available at https://anonymous.4open.science/r/SCAM-BDD3.
Machine Learning, ICML
@@ 1 Introduction
Time Series Forecasting (TSF) is a crucial task with extensive applications in energy, finance, engineering, and many other domains. Recent advances in deep learning have resulted in TSF methods that outperform traditional methods in precision, robustness, and scalability (Zhou et al., 2021; Salinas et al., 2020; Taylor & Letham, 2018).
Nevertheless, deep learning-based TSF methods still face significant challenges such as overfitting, dependence on high-quality datasets, and inconsistent performance across datasets — issues exacerbated by flawed evaluation practices (Shao et al., 2024; Qiu et al., 2024). Central to these challenges is the problem of low-quality labels [^1] In TSF, given two adjacent windows in a time series, $x$ (input window) and $y$ (output window), “labels” refer to $y$ fitted by predictions $\hat{y}=f(x;\theta)$ in a supervised learning setting., associated with inherent noise and anomalies in raw data. Existing strategies, such as multimodal data integration (Xu et al., 2024a; Chen et al., 2024a) and dataset scaling (Shi et al., 2024), focus on augmenting or refining datasets but fail to address the core limitation: the reliance on raw labels as ground truth. To address this, we pose two critical questions:
@@
(a) Grid search
(b) Co-optimization
Figure 1: Illustrations of the proposed method Scam.
Specifically, our approach optimizes a simple reconstruction network $g(\cdot;\phi)$ to generate intermediate reconstructions. Each parameter $\phi_{i}$ during optimization corresponds to an individual candidate dataset $D_{i}$ (see Figure 1(a)). Labels are collected from each candidate dataset to train an individual predictor $f(\cdot;\theta)$ under supervised settings. The two-step hierarchical optimization functions as a search on an auxiliary reconstruction metric (see Section 3.1). To integrate this process into a feasible pipeline, we simplify this process as a one-step optimization for a co-objective (see Section 3.2).
The optimal generalized performance lies on the Pareto Front of the co-objective optimization. However, a critical issue is overfitting, as depicted in Figure 1(b), which prevents convergence to optimality when the trajectory moves to one extreme of the Pareto Front (where the reconstruction closely approximates the raw data). To better analyze overfitting, we derive a mask form of the co-objective loss, which partitions original time series into distinct components. We employ a criterion, namely the sharpness metric $\lambda_{max}$ (Ilbert et al., 2024), which detects overfitting by evaluating the sharpness of loss landscape. Using $\lambda_{max}$ and practical evaluations of the decomposed loss, we identify the overfitted components. This process, termed Self-Correction with Adaptive Mask (Scam), discards raw data labels ($Y$) based on reconstruction results ($\tilde{Y}$) and current predictions ($\hat{Y}$), smoothing the loss landscape and enhancing the generalization of the predictor model $f(\cdot;\theta)$.
[…]
Figure 2: Working Pipeline of Scam. $g(\cdot;\phi)$ is a reconstruction network; $f(\cdot;\theta)$ is a predictor; $\mathcal{L}$ uses reconstructed labels $\tilde{Y}$ to mask overfitted components.
Figure 2 depicts the general working pipeline, where we apply masked updates on a reconstruction network $g(\cdot;\phi)$ and a predictor model $f(\cdot;\theta)$ during training. During inference, the prediction will be generated directly by the updated $f(\cdot;\theta)$ with no additional cost. To further generalize across models of varying complexities, we also propose to use Spectral Norm Regularization (Yoshida & Miyato, 2017; Miyato et al., 2018).
@@ 2.1 Problem Formulation
[…]
@@
$D_{i}$ =$=\left\{\tilde{X}_{i},\tilde{Y}_{i}\right\}$ (1)
=$=\left\{\left\{g\left(x;\phi_{i}\right)\mid x\in X_{i}\right\},\left\{g\left(y;\phi_{i}\right)\mid y\in Y_{i}\right\}\right\},$
@@
$\|f(X_{test};\theta^{\ast}(\phi^{\ast}))-Y_{test}\|\leq\|f(X_{test};\theta^{\ast}(\phi^{i}))-Y_{test}\|,\forall\phi_{i},$
where $\theta^{\ast}(\phi^{i})$ indicates that $\theta^{\ast}$ is optimized on $D_{trn}^{i}=\{\tilde{X}_{i},\tilde{Y}_{i}\}=\{\{g(x_{j};\phi_{i})\},\{g(y_{j};\phi_{i})\}\}$.
@@ 2.2 Proposed $g(\cdot;\phi)$ for Reconstruction
As depicted in Figure 3, the proposed $g(\cdot;\phi)$ comprises four convolutional layers, a cross-layer concatenation for multi-resolution integration, and a lightweight feedforward network (FFN) to decode the reconstructed results. The convolutional layers primarily act as a parameterized smoothing mechanism, similar to techniques for seasonal-trend decomposition (Zeng et al., 2023; Lin et al., 2024a, b). The FFN then mixes information from different positions in a time series to reconstruct each data point using features extracted by convolutional layers (further details in Appendix C). Without introducing extra noise or patch-wise/point-wise masks (Nie et al., 2022), we directly learn an identity mapping.
In essence, $g(\cdot;\phi)$ is designed to traverse diverse alternatives with varying levels of fidelity, which differs in purpose from the usual reconstruction task (Liu et al., 2024c; Ma et al., 2022; Liu & Chen, 2024). Due to this difference, the architecture of $g(\cdot;\phi)$ might benefit from novel designs; however, this is not examined in this paper. The proposed $g(\cdot;\phi)$ is merely a simple prototype to validate our ideas and is not claimed superior to other unexplored options for this novel task.
@@ 3.1 Initial Case
In this setup, we optimize $g(\cdot;\phi)$ for reconstruction loss using full-batch gradient descent, while the predictor $f(\cdot;\theta)$ is optimized with mini-batch SGD in a vanilla training setting. At each optimization epoch $i$, we freeze the current parameters of $g(\cdot;\phi_{i})$ and generate a candidate dataset $D_{i}$. The predictor $f(\cdot;\theta)$ is re-initialized before the epoch and trained on $D_{i}$ until convergence. The prediction results on the original test set are recorded for each $g(\cdot;\phi_{i})$ and corresponding trained $f(\cdot;\theta^{\ast}_{i})$. A case experiment is conducted on the ETTh1 dataset using a vanilla 2-layer MLP model (for faster and more guaranteed convergence). The pseudo-code for this grid search is provided in Appendix A.
@@
Figure 4: Grid search results. Each scatter denotes a converged predictor $f(\cdot;\theta)$ on an individual dataset parameterized by $g(\cdot;\phi_{i})$.
@@
Figure 4(a) demonstrates the improved performance of better-labeled candidates, as indicated by the scatter points below the red dotted line — which represents the baseline performance of the predictor trained on the raw dataset.
@@ 3.2 Co-objective Training
A single-step optimization using mini-batch SGD would be sufficient, enabling a more smooth trajectory of $\phi$ during updates (Figure 1(b) vs 1(a) and Figure 5(a) vs Figure 4(b)). Moreover, enabling gradient computation of $\ell_{pred}=\|\tilde{y}-\hat{y}\|$ w.r.t. $\phi$ introduces a regularization effect, making $\tilde{y}$ updated more cautiously towards $y$. This allows us to constrain the update of candidate datasets (now specifically represented by the reconstructed $\tilde{y}$) by jointly constraining gradients w.r.t. both $\theta$ and $\phi$:
@@
The gradient constraint has a surrogate $\|\nabla_{\theta}f\|$ and practically implemented by Spectral Norm Regularization (snr), as discussed in Section 3.4 and Appendix B. For simplicity, we temporarily omit this constraint, as a 2-layer MLP converges quickly with a small $\|\nabla_{\theta}f\|$. Using the same setting as the grid search, we evaluate the revised loss function for co-optimizing the predictor and the reconstruction network. Figure 5 shows that this co-training improves $\ell_{target}$ loss while simplifying the two-step training process, leading to more stable optimization and reduced training costs.
@@
(a) Loss curve of co-objective
(b) Simplified Pareto Front
Figure 5: Loss curve and Pareto Front of co-objective training.
@@ 3.3 Self-Correction with Adaptive Mask (Scam)
Figure 6: Loss curves of 4 parts divided by the adaptive mask. From left to right, $2|\tilde{y}-y|\odot\overline{M_{<}}\odot M$, $2|\tilde{y}-\hat{y}|\odot M_{<}\odot M$, $|y-\hat{y}|\odot M$ and $|y-\hat{y}|\odot\overline{M}$. The four parts combined are the loss $\mathcal{L}$ in Eq. 5. $M$ refers to the mask where $m_{i}>0$ and $\overline{M}$ refers to the mask where $m_{i}<0$. The right y-axis (in green) indicates the value for $(\ell_{test}-\ell_{train})$.
Mask Form of Self-supervised Loss. In a traditional supervised-learning paradigm, the target loss $\ell_{target}=\|\hat{y}-y\|$ is used only to train a model, implying that all data points are equally treated as valid labels. However, in our approach, where predictors are trained alongside a search for candidate datasets guided by the reconstruction loss, the labels perceived by the predictors are adaptively shifted. Specifically, $\hat{y}=f(x;\theta)$ is trained to fit $\tilde{y}$, meaning only the second term is optimized for the predictor $f(\cdot;\theta)$ (Eq. 3). The reconstruction loss term, on the other hand, is optimized to provide self-supervised labels for the predictor. We frame this as a self-supervised-learning paradigm that adaptively adjusts labels in TSF problems. By comparing the revised loss with the traditional supervised loss, we can explicitly separate the auxiliary loss from the supervised loss:
$\mathcal{L}$ $=\underbrace{\left(y-\hat{y}\right)^{2}}_{\mathcal{L}_{sup}}+\underbrace{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\Big{[}\left(\tilde{y}-y\right)^{2}+\left(\hat{y}-\tilde{y}\right)^{2}-\left(y-\hat{y}\right)^{2}\Big{]}}}_{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\mathcal{L}_{aux}}}$ (4)
$=\mathcal{L}_{sup}+{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}2\left(\tilde{y}-\hat{y}\right)\left(\tilde{y}-y\right)}.$
When revisiting the objective $\mathcal{L}$ in co-training, the additional loss term $\mathcal{L}_{aux}$ does not directly contribute to the target objective. Instead, this term depends on the relative positions of $\hat{y}$, $\tilde{y}$ and $y$. When the reconstructed $\tilde{y}$ is viewed as a correction of labels, $\mathcal{L}_{aux}$ indicates where the correction should be placed. Time series are naturally sparse in real scenarios, often containing spiking signals due to irregular events or anomalies. $\mathcal{L}_{aux}$ encourages the reconstructed $\tilde{y}$ to lie between the prediction $\hat{y}$ and the actual labels $y$, which can undermine sparsity when used as labels.
Eq. 4 is based on $\ell_{2}$-norm (Mean Squared Error, MSE). Alternatively, we can adopt the more error-robust $\ell_{1}$-norm (Mean Absolute Error, MAE) to reformulate:
$\mathcal{L}$ $=|y-\hat{y}|+{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\left(|\tilde{y}-\hat{y}|+|\tilde{y}-y|-|y-\hat{y}|\right)}$ (5)
@@
Let $\text{Let }m=(\tilde{y}-\hat{y})(\tilde{y}-y),$
@@
Here, $M$ is a binary mask defined by $m=(\tilde{y}-\hat{y})(\tilde{y}-y)>0$, $M_{<}$ ensures $|\tilde{y}-\hat{y}|<|\tilde{y}-y|$, and $\overline{M_{<}}$ represents its complement. $M$ functions similarly to $\mathcal{L}_{aux}$ in Eq. 4 while $M_{<}$ ensures $\tilde{y}$ is optimized with a relatively small margin.
@@
Further evidence is provided by analyzing the loss landscape. A sharpness metric for optimization, proposed by (Ilbert et al., 2024), measures the generalization capability of a model. Specifically, the sharpness metric is defined as $\lambda_{max}=\|\nabla^{2}_{\theta}\mathcal{L}\|_{2}^{2}$, the largest eigenvalue of the Hessian matrix. A higher $\lambda_{max}$ indicates a sharper loss landscape, which correlates with a worse generalization or more severe overfitting. By computing this metric on the converged parameters, we observe that $\ell_{target}\odot M$ does exhibit a sharper landscape compared to $\ell_{target}\odot\overline{M}$. To address this, we keep the less sharp part of the $\mathcal{L}_{sup}$ term, i.e., $\mathcal{L}_{sup}\odot\overline{M}$, resulting in the revised loss:
$\mathcal{L}$ $=\|y-\hat{y}\|{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\ \odot\ \overline{M}}$ (6)
$+{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}2\left(\|\tilde{y}-\hat{y}\|\odot M_{<}+\|\tilde{y}-y\|\odot\overline{M_{<}}\right)\odot M}.$

### v3

@@ Abstract
Time Series Forecasting (TSF) is a crucial task in various domains, yet existing TSF models rely heavily on high-quality data and insufficiently exploit all available data. This paper explores a novel self-supervised approach to re-label time series datasets by inherently constructing candidate datasets. During the optimization of a simple reconstruction network, intermediates are used as pseudo labels in a self-supervised paradigm, improving generalization for any predictor. We introduce the Self-Correction with Adaptive Mask (Scam), which discards overfitted components and selectively replaces them with pseudo labels generated from reconstructions. Additionally, we incorporate Spectral Norm Regularization (snr) to further suppress overfitting from a loss landscape perspective. Our experiments on eleven real-world datasets demonstrate that Scam consistently improves the performance of various backbone models. This work offers a new perspective on constructing datasets and enhancing the generalization of TSF models through self-supervised learning. The code is available at https://github.com/SuDIS-ZJU/SCAM.
@@ 1 Introduction
Time Series Forecasting (TSF) is a crucial task with extensive applications in energy, finance, engineering, and many other domains. Recent advances in deep learning have resulted in TSF methods that outperform traditional methods in precision, robustness, and scalability [53, 31, 36].
Nevertheless, deep learning-based TSF methods still face significant challenges such as overfitting, dependence on high-quality datasets, and inconsistent performance across datasets — issues exacerbated by flawed evaluation practices [34, 30]. Central to these challenges is the problem of low-quality labels [^1] In TSF, given two adjacent windows in a time series, $x$ (input window) and $y$ (output window), “labels” refer to $y$ fitted by predictions $\hat{y}=f(x;\theta)$ in a supervised learning setting., associated with inherent noise and anomalies in raw data. Existing strategies, such as multimodal data integration [43, 4] and dataset scaling [35], focus on augmenting or refining datasets but fail to address the core limitation: the reliance on raw labels as ground truth. To address this, we pose two critical questions:
@@
Figure 1: The illustration of the proposed method Scam.
From an intuitive perspective, a straightforward reconstruction network $g(\cdot;\phi)$ can be employed to generate pseudo-labels in a self-supervised manner. To assess the feasibility of training such a reconstruction network, we conduct an initial evaluation using a grid search algorithm (Section 3.1, Appendix B). The network $g$ is optimized to learn an identity mapping (i.e., reconstruction) of the raw labels over multiple training epochs. During each epoch, the intermediate parameters of $g$ can be frozen to produce an approximate reconstruction of the dataset. Empirical observations reveal that when certain predictors $f(\cdot;\theta)$ are trained from scratch on these intermediate datasets, they exhibit superior performance compared to their counterparts trained exclusively on the raw datasets.
While the reconstruction-based approach demonstrates potential for pseudo-label generation, determining optimal parameters $\phi$ presents a significant challenge. Specifically, the reconstruction loss (i.e., $\|y-\tilde{y}\|$), when combined with the original supervised loss (i.e., $\|\hat{y}-y\|$), tends to drive $g(\cdot;\phi)$ toward reproducing the raw labels. This phenomenon induces an undesirable "overfitting" behavior during training.
To better analyze overfitting, we derive a mask form of the co-objective loss, which partitions original time series into distinct components (see Figure 1). We employ a criterion, namely the sharpness metric $\lambda_{max}$ [10], which detects overfitting by evaluating the sharpness of loss landscape. Using $\lambda_{max}$ and practical evaluations of the decomposed loss, we identify the overfitted components. This process, termed Self-Correction with Adaptive Mask (Scam), discards raw data labels ($y$) based on reconstruction results ($\tilde{y}$) and current predictions ($\hat{y}$), smoothing the loss landscape and enhancing the generalization of the predictor model $f(\cdot;\theta)$.
We apply masked updates on a reconstruction network $g(\cdot;\phi)$ and a predictor model $f(\cdot;\theta)$ during training. During inference, the prediction will be generated directly by the updated $f(\cdot;\theta)$ with no additional cost. To further generalize across models of varying complexities, we also propose to apply Spectral Norm Regularization [46, 26] to parameters of the first or last linear layers without altering optimizers like previous works [10].
@@ 2.1 Problem Formulation
A typical TSF process is formulated as a supervised-learning problem, i.e., to find $\theta^{\ast}=\arg\underset{\theta}{\min}\left\|f(X;\theta)-Y\right\|$, where a specified metric $\|\cdot\|$ is used to measure errors, typically the $\ell_{1}$- or $\ell_{2}$-norm.
@@
$D_{i}$ $=\left\{\tilde{X}_{i},\tilde{Y}_{i}\right\}=\left\{\left\{g\left(x;\phi_{i}\right)\mid x\in X_{i}\right\},\left\{g\left(y;\phi_{i}\right)\mid y\in Y_{i}\right\}\right\},$ (1)
@@
$\|f(X_{test};\theta^{\ast}(\phi^{\ast}))-Y_{test}\|\leq\|f(X_{test};\theta^{\ast}(\phi_{i}))-Y_{test}\|,\forall\phi_{i},$
where $\theta^{\ast}(\phi_{i})$ indicates that $\theta^{\ast}$ is optimized on $D_{trn}^{i}=\{\tilde{X}_{i},\tilde{Y}_{i}\}=\{\{g(x_{j};\phi_{i})\},\{g(y_{j};\phi_{i})\}\}$.
@@ 2.2 Proposed $g(\cdot;\phi)$ for Reconstruction
As depicted in Figure 2, the proposed $g(\cdot;\phi)$ comprises four convolutional layers, a cross-layer concatenation for multi-resolution integration, and a lightweight feedforward network (FFN) to decode the reconstructed results. The convolutional layers primarily act as a parameterized smoothing mechanism, similar to techniques for seasonal-trend decomposition [48, 16, 17]. The FFN then mixes information from different positions in a time series to reconstruct each data point using features extracted by convolutional layers (further details in Appendix D). Without introducing extra noise or patch-wise/point-wise masks [27], we directly learn an identity mapping.
In essence, $g(\cdot;\phi)$ is designed to traverse diverse alternatives with varying levels of fidelity, which differs in purpose from the usual reconstruction task [22, 24, 18]. Due to this difference, the architecture of $g(\cdot;\phi)$ might benefit from novel designs; however, this is not examined in this study. The proposed $g(\cdot;\phi)$ is merely a simple prototype to validate our ideas and is not claimed superior to other unexplored options for this novel task.
@@ 3.1 Initial Case
In this setup, we optimize $g(\cdot;\phi)$ for reconstruction loss using full-batch gradient descent, while the predictor $f(\cdot;\theta)$ is optimized with mini-batch SGD in a vanilla training setting. At each optimization epoch $i$, we freeze the current parameters of $g(\cdot;\phi_{i})$ and generate a candidate dataset $D_{i}$. The predictor $f(\cdot;\theta)$ is re-initialized before the epoch and trained on $D_{i}$ until convergence. The prediction results on the original test set are recorded for each $g(\cdot;\phi_{i})$ and corresponding trained $f(\cdot;\theta^{\ast}_{i})$. A case experiment is conducted on the ETTh1 dataset using a vanilla 2-layer MLP model (for faster and more guaranteed convergence). The pseudo-code for this grid search is provided in Appendix B.
@@
Figure 3: (a) shows the distribution of candidate datasets during the grid search. (b) shows the unsmooth loss curve in grid search. Each scatter in both figures denotes a converged predictor $f(\cdot;\theta)$ on an individual dataset parameterized by $g(\cdot;\phi_{i})$.
@@
Figure 3(a) demonstrates the existence of better-labeled candidates which bring better performance for a given predictor, as indicated by the scatter points below the red dotted line — which represents the baseline performance of the predictor trained on the raw dataset.
@@ 3.2 Co-objective Training
(a) Loss curve of co-objective
(b) Simplified Pareto Front
Figure 4: (a) shows the overfitting tendency supported by the increasing $\ell_{target}$ which measures errors on the test set. (b) shows the Pareto Front of two objectives, tilting top-left indicating the overfitting in (a).
@@
A single-step optimization using mini-batch SGD would be sufficient, enabling a more smooth trajectory of $\phi$ during updates (Figure 5(b) vs 5(a) and Figure 4(a) vs Figure 3(b)). Moreover, enabling gradient computation of $\ell_{pred}=\|\tilde{y}-\hat{y}\|$ w.r.t. $\phi$ introduces a regularization effect, making $\tilde{y}$ updated more cautiously towards $y$. This allows us to constrain the update of candidate datasets (now specifically represented by the reconstructed $\tilde{y}$) by jointly constraining gradients w.r.t. both $\theta$ and $\phi$:
@@
$\|\nabla_{\theta,\phi}\tilde{y}_{i}\|\leq\delta,\quad\forall\tilde{y}_{i}\in\tilde{Y}$
@@
(a) Grid search
(b) Co-optimization
Figure 5: Illustrations of grid search and co-optimization.
The gradient constraint has a surrogate $\|\nabla_{\theta}f\|$ and practically implemented by Spectral Norm Regularization (snr), as discussed in Section 3.4 and Appendix C. For simplicity, we temporarily omit this constraint, as a 2-layer MLP converges quickly with a small $\|\nabla_{\theta}f\|$. Using the same setting as the grid search, we evaluate the revised loss function for co-optimizing the predictor and the reconstruction network. Figure 4 shows that this co-training improves $\ell_{target}$ loss while simplifying the two-step training process, leading to more stable optimization and reduced training costs.
@@ 3.3 Self-Correction with Adaptive Mask (Scam)
Figure 6: Loss curves of 4 parts divided by the adaptive mask. From left to right, $2|\tilde{y}-y|\odot\overline{M_{<}}\odot M$, $2|\tilde{y}-\hat{y}|\odot M_{<}\odot M$, $|y-\hat{y}|\odot M$ and $|y-\hat{y}|\odot\overline{M}$. The four parts combined are the loss $\mathcal{L}$ in Eq. 4. $M$ refers to the mask where $m_{i}>0$ and $\overline{M}$ refers to the mask where $m_{i}<0$. The right y-axis (in green) indicates the value for $(\ell_{test}-\ell_{train})$, the differences in losses measured on the test set and the train set. Values (on right y-axis) above 0 (the black dotted line) indicate overfitting.
Mask Form of Self-supervised Loss. In a traditional supervised-learning paradigm, the target loss $\ell_{target}=\|\hat{y}-y\|$ is used only to train a model, implying that all data points are equally treated as valid labels. However, in our approach, where predictors are trained alongside a search for candidate datasets guided by the reconstruction loss, the labels perceived by the predictors are adaptively shifted. Specifically, $\hat{y}=f(x;\theta)$ is trained to fit $\tilde{y}$, meaning only the second term is optimized for the predictor $f(\cdot;\theta)$ (Eq. 3). The reconstruction loss term, on the other hand, is optimized to provide self-supervised labels for the predictor. We frame this as a self-supervised-learning paradigm that adaptively adjusts labels in TSF problems. By comparing the revised loss with the traditional supervised loss, we can explicitly separate the auxiliary loss from the supervised loss: When revisiting the objective $\mathcal{L}$ in co-training, the additional loss term $\mathcal{L}_{aux}$ does not directly contribute to the target objective. Instead, this term depends on the relative positions of $\hat{y}$, $\tilde{y}$ and $y$. When the reconstructed $\tilde{y}$ is viewed as a correction of labels, $\mathcal{L}_{aux}$ indicates where the correction should be placed. Time series are naturally sparse in real scenarios, often containing spiking signals due to irregular events or anomalies. $\mathcal{L}_{aux}$ encourages the reconstructed $\tilde{y}$ to lie between the prediction $\hat{y}$ and the actual labels $y$, which can undermine sparsity when used as labels.
$\mathcal{L}$ $=\underbrace{|y-\hat{y}|}_{\mathcal{L}_{sup}}+\underbrace{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\left(|\tilde{y}-\hat{y}|+|\tilde{y}-y|-|y-\hat{y}|\right)}}_{\mathcal{L}_{aux}}$ (4)
@@
$\text{Let }m=(\tilde{y}-\hat{y})(\tilde{y}-y),$
@@
Here, $M$ is a binary mask defined by $m=(\tilde{y}-\hat{y})(\tilde{y}-y)>0$, $M_{<}$ ensures $|\tilde{y}-\hat{y}|<|\tilde{y}-y|$, and $\overline{M_{<}}$ represents its complement.
@@
Further evidence is provided by analyzing the loss landscape. A sharpness metric for optimization, proposed by [10], measures the generalization capability of a model. Specifically, the sharpness metric is defined as $\lambda_{max}=\|\nabla^{2}_{\theta}\mathcal{L}\|_{2}^{2}$, the largest eigenvalue of the Hessian matrix. A higher $\lambda_{max}$ indicates a sharper loss landscape, which correlates with a worse generalization or more severe overfitting. By computing this metric on the converged parameters, we observe that $\ell_{target}\odot M$ does exhibit a sharper landscape compared to $\ell_{target}\odot\overline{M}$. To address this, we keep the less sharp part of the $\mathcal{L}_{sup}$ term, i.e., $\mathcal{L}_{sup}\odot\overline{M}$, resulting in the revised loss:
$\mathcal{L}=\underbrace{\|y-\hat{y}\|{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\ \odot\ \overline{M}}}_{\text{masked $\mathcal{L}_{sup}$}}+\underbrace{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}2\left(\|\tilde{y}-\hat{y}\|\odot M_{<}+\|\tilde{y}-y\|\odot\overline{M_{<}}\right)\odot M}}_{\mathcal{L}_{aux}}.$ (5)
@@
3.4 Spectral Norm Regularization (snr)
Note that we have omitted the gradient constraints in Eq. 3. This term is, in fact, positively correlated with $\nabla_{\theta}f(x;\theta)$ (discussed in Appendix E). This further supports our previous analysis, as we have only tested MLP models, which converge easily with lower $\nabla_{\theta}f$. However, when the predictor is replaced by a Transformer-based model, the dominant source of overfitting shifts from $\ell_{target}$ to $\ell_{pred}$. This phenomenon is not unique to our self-supervised learning paradigm. For example, [10] proposed sharpness-aware optimization to address overfitting in supervised settings. While gradient penalties are theoretically effective, they may not be practical for complex models like Transformers, as computing second-order derivatives can significantly hinder optimization.